A 750 kg car slows to a stop from traveling southbound at 30 m/s over a distance of 500 m. What was the acceleration needed to stop the car? How large was the force required to stop the car?

Two cables are being used to pull a 100 kg sled on a frictionless surface. The first cable has a tension of 500 N in a direction 15o south of east. The other cable has a tension of 800 N in a direction 30o east of north. What is the total horizontal force on the sled?

A 50 kg crate is being lowered by a cable. (a) If it is held still, what is the tension in the cable? (b) If it is being lowered with an acceleration of 1.0 m/s2, what is the tension in the cable? (c) If it is being raised with an velocity of 1.0 m/s , what is the tension in the cable?
Answer #1

A)

mass of car,m- 750kg

initial velocity ,30m/s

final velocity ,u = Orns /

distance to stop,s500m

a=?

Use formula v^{2}-u^{2}=2as

(Om/s)- (30m/s)2a500m

02-302 2 * a * 500

-302 / 1000-а

ANSWER: {color{Red} a=-0.9m/s^{2}}

----------------------------

Force=?

Use Formula F=ma

F 750kg * 0.9m/s

ANSWER: F 675

=====================================

B)

800N 30deg 500N 15 deg

Horizontal Force,F_{x}=500Ncos15 +800Ncos60

F_{x}=500cos15 +800cos60

ANSWER: {color{Red} F_{x}=882.96N}

=======================

C)

(a)Held still

Tension= weight of crate =mg =50kg*9.81m/s2

ANSWER: {color{Red} T=490.5N}

(b) If it is being lowered with an acceleration of 1.0 m/s2,

T=mg-ma

T=m(g-a)

T=50kg(9.81m/s^{2}-1m/s^{2})

T=50*8.81

ANSWER:{color{Red} T=440.5N}

====================

c)If it is being raised with an velocity of 1.0 m/s

Use Formula T=mg+ma

but since velocity is constant, acceleration is zero

T=mg+m*0

T=mg

T = 50kg * 9.81m/

ANSWER: {color{Red} T=490.5N}

=====================